# High-Performance Data Format for Scientific Data Storage and Analysis

### v1

@@ Abstract
In this article, we present the High-Performance Output (HiPO) data format developed at Jefferson Laboratory for storing and analyzing data from Nuclear Physics experiments. The format was designed to efficiently store large amounts of experimental data, utilizing modern fast compression algorithms. The purpose of this development was to provide organized data in the output, facilitating access to relevant information within the large data files. The HiPO data format has features that are suited for storing raw detector data, reconstruction data, and the final physics analysis data efficiently, eliminating the need to do data conversions through the lifecycle of experimental data. The HiPO data format is implemented in C++ and JAVA, and provides bindings to FORTRAN, Python, and Julia, providing users with the choice of data analysis frameworks to use. In this paper, we will present the general design and functionalities of the HiPO library and compare the performance of the library with more established data formats used in data analysis in High Energy and Nuclear Physics (such as ROOT and Parquete).
@@ 4.1 Design
Listing 5: c++ example to read HiPO tuple file and fill a histogram.
@@
The example code in Listing 5 demonstrates how to open a file and read the branches named ”c1” through ”c4.” It processes all rows in the file, passing them to the user code via a declared array. The values from the first branch (”c1”) are then used to populate a histogram.
@@ 4.2 Benchmarking
The numerical results of these tests are summarized in Table 1. The HiPO data format demonstrates significantly faster reading speeds compared to Parquet and ROOT (RNTuple) when performing simple data reading in a loop (as illustrated in Listing 5), in both C++ and Java implementations. The ”DataFrame” column in the table represents an experimental HiPO feature where histograms are passed directly to the tuple reader, utilizing bulk histogram filling from data buckets. This method achieves a $50\%$ improvement in histogram filling efficiency.
@@
Listing 6: Simplified version for filling histograms to compare data reading speeds more accurately.
During our tests, we observed a performance difference between TH1D.Fill() in ROOT and twig::h1d.fill() (our custom implementation of a fast histogramming package), which impacted the overall benchmark results. To eliminate this variable, we replaced the histogramming code in both cases with a simplified version, as shown in Listing 6, and reran the benchmarks. The execution times obtained were $1.589$ seconds for ROOT (RNTuple) and $0.523$ seconds for HiPO, demonstrating that HiPO achieves read-and-fill times three times faster than ROOT RNTuple.

### v3

@@ Abstract
In this article, we present the High-Performance Output (HiPO) data format developed at Jefferson Laboratory for storing and analyzing data from Nuclear Physics experiments. The format was designed to efficiently store large amounts of experimental data, utilizing modern fast compression algorithms. The purpose of this development was to provide organized data in the output, facilitating access to relevant information within the large data files. The HiPO data format has features that are suited for storing raw detector data, reconstruction data, and the final physics analysis data efficiently, eliminating the need to do data conversions through the lifecycle of experimental data. The HiPO data format is implemented in C++ and JAVA, and provides bindings to FORTRAN, Python, and Julia, providing users with the choice of data analysis frameworks to use. In this paper, we will present the general design and functionalities of the HiPO library and compare the performance of the library with more established data formats used in data analysis in High Energy and Nuclear Physics (such as ROOT and Parquete). In columnar data analysis, HiPO surpasses established data formats in performance and can be effectively applied to data analysis in other scientific fields.
@@ 17 Bank b = new Bank(schema,2); // create a table with 2 rows
4 Event Tagging
The HiPO is designed to store data in units called events, which are collections of related data, typically organized by time. In nuclear physics experiments, an event corresponds to a single interaction between a beam particle and a target, capturing the resulting particles from that interaction. After data acquisition, post-processing identifies the number and properties of the particles produced, recording this information within the event for further physics analysis. Different interactions occur at varying frequencies in these experiments, and specific analysis programs focus on particular physics reactions. These programs do not analyze every event but instead require events with a certain number of reconstructed particles to proceed.
Figure 2 illustrates a schematic view of how data can be theoretically distributed within a file. Event types are represented by rows in a table listing the particles reconstructed by data processing software. For rare interactions, much of the data read is irrelevant to the analysis, leading to inefficient use of I/O resources. This inefficiency places unnecessary strain on shared file systems and computational resources.
Figure 2: Schematic view of the file output for tagged and untagged events. When using the tagged mode of writing a file, events are organized in the records based on the tag of the event, which is assigned by the user depending on the needs.
To address this issue, the HiPO data format incorporates a tagging feature for organizing stored data. Each record in the file is assigned a unique tag identifier, and this information is stored in the file footer along with the record’s positions and sizes. This approach enables events to be sorted into separate records during the file-writing process, grouping similar events together. Listing 5 provides an example of tagging events based on the number of rows in a table that lists clusters, as described in the previous example.
1// Writing arrays into an Event
2HipoReader r = new HipoReader("myfile.h5");
3// using HipoWriter.create() transfers all the dictionary
4// and the metadata to the write object
5HipoWriter w = HipoWriter.create("taggedfile.h5",r);
6
7Bank[] b = r.getBanks("data::clusters");
8Event event = new Event();
9while(r.next(event)){
10 event.read(b);
11 if(b[0].getRows()>=5) event.setTag(5);
12 if(b[0].getRows()==2) event.setTag(2);
13 if(b[0].getRows()==3) event.setTag(3);
14 w.addEvent(event);
15}
16w.close();
Listing 5: Java example to create and write primitive types into an event
The example code generates a file in which all events with the same number of rows are grouped together. Events with a specific number of rows in the ”clusters” bank can be accessed directly, eliminating the overhead of scanning through the entire file. An example of reading events where the number of rows in the ”clusters” bank is equal to 2 or greater than 4 is provided in Listing 6.
1// Writing arrays into an Event
2HipoReader r = new HipoReader();
3r.setTags(2,5); // read only tag=2 and tag=5
4r.open("taggedfile.h5");
5Bank[] b = r.getBanks("data::clusters");
6while(r.nextEvent(b)){
7 b[0].show();// print bank content on the screen
8}
Listing 6: Java example to sort events in the output file depending on number of rows in the table
In Listing 5, only the ”data::clusters” bank containing either two rows or more than four rows is read. However, when an event is written to the output file, the entire event, including all banks and nodes, is recorded. The decision on how to partition the file, though, is based solely on the contents of the specified bank. This feature is utilized in CLAS12 to tag events based on their topology, such as the number and types of detected particles. This approach simplifies the process of reading only the relevant events from the data file for analysis.
@@ 5.1 Design
Listing 7: c++ example to read HiPO tuple file and fill a histogram.
@@
The example code in Listing 7 demonstrates how to open a file and read the branches named ”c1” through ”c4.” It processes all rows in the file, passing them to the user code via a declared array. The values from the first branch (”c1”) are then used to populate a histogram.
@@ 5.2 Benchmarking
The numerical results of these tests are summarized in Table 1. The HiPO data format demonstrates significantly faster reading speeds compared to Parquet and ROOT (RNTuple) when performing simple data reading in a loop (as illustrated in Listing 7), in both C++ and Java implementations. The ”DataFrame” column in the table represents an experimental HiPO feature where histograms are passed directly to the tuple reader, utilizing bulk histogram filling from data buckets. This method achieves a $50\%$ improvement in histogram filling efficiency.
@@
Listing 8: Simplified version for filling histograms to compare data reading speeds more accurately.
During our tests, we observed a performance difference between TH1D.Fill() in ROOT and twig::h1d.fill() (our custom implementation of a fast histogramming package), which impacted the overall benchmark results. To eliminate this variable, we replaced the histogramming code in both cases with a simplified version, as shown in Listing 8, and reran the benchmarks. The execution times obtained were $1.589$ seconds for ROOT (RNTuple) and $0.523$ seconds for HiPO, demonstrating that HiPO achieves read-and-fill times three times faster than ROOT RNTuple.